# ManipVQA: Injecting Robotic Affordance and Physically Grounded Information into Multi-Modal Large Language Models

### v1

@@ Abstract
The integration of Multimodal Large Language Models (MLLMs) with robotic systems has significantly enhanced the ability of robots to interpret and act upon natural language instructions. Despite these advancements, conventional MLLMs are typically trained on generic image-text pairs, lacking essential robotics knowledge such as affordances and physical knowledge, which hampers their efficacy in manipulation tasks. To bridge this gap, we introduce ManipVQA, a novel framework designed to endow MLLMs with Manipulation-centric knowledge through a Visual Question-Answering format. This approach not only encompasses tool detection and affordance recognition but also extends to a comprehensive understanding of physical concepts. Our approach starts with collecting a varied set of images displaying interactive objects, which presents a broad range of challenges in tool object detection, affordance, and physical concept predictions. To seamlessly integrate this robotic-specific knowledge with the inherent vision-reasoning capabilities of MLLMs, we adopt a unified VQA format and devise a fine-tuning strategy that preserves the original vision-reasoning abilities while incorporating the new robotic insights. Empirical evaluations conducted in robotic simulators and across various vision task benchmarks demonstrate the robust performance of ManipVQA. Code and dataset will be made publicly available at https://github.com/SiyuanHuang95/ManipVQA.
@@ I Introduction
Recently, Multi-modality Large Language Models such as OpenAI’s GPT-4(V) [1], Google’s Gemini and SPHINX-X [2] have made significant strides in enhancing the ability to perceive and comprehend both text and images. These models typically aim to align multi-modal encoders with large language models, a feat achieved by training on numerous text-image pairs or interleaved text-image pairs to foster a more comprehensive understanding of both modalities. These large-scale vision-language models often exhibit promising potential in addressing common sense reasoning and demonstrate remarkable generalization in vision tasks. However, the application of MLLMs in manipulation tasks [3, 4], such as robotic affordance and the design of such robotic systems, remain open challenges.
@@
Robotic manipulation [5, 4] is defined as a robot’s capacity to perceive its environment and discern the potential actions applicable to various objects. The direct application of existing MLLMs, pretrained on common scene reasoning, to robotic manipulation, does not yield satisfactory performance due to the absence of low-level action samples in their pretraining data. Previous research has addressed robotic affordance by prompting MLLMs to process scene images and subsequently generate a sequence of robotic actions, thus not achieving promising performance. Affordance grounding [6] aims to localize the regions in objects where actions are possible. This task also faces the challenge of establishing an explicit link with object parts due to the diversity of interactive affordances. Prior research [7] employs Large Language Models (LLMs) to get the answer on affordance with prompt engineering. Despite the potential utility of MLLMs in robotic applications, their effectiveness is curtailed by several challenges. Classic MLLMs [2] are generally trained on generic image-text pairs, thereby lacking critical robotic knowledge such as affordances and physical principles. This deficiency in specialized knowledge impedes their performance in manipulation tasks, consequently restricting the variety of tasks that robots can execute and the precision of their execution. The discrepancy between the capabilities of MLLMs and the demands of robotic systems constitutes a significant challenge that needs to be addressed.
@@
To address this gap, we present ManipVQA, an innovative framework devised to equip MLLMs with manipulation-centric knowledge via a Visual Question-Answering paradigm. This integration is achieved through a unified VQA approach and a fine-tuning strategy that preserves the original vision-reasoning strengths of MLLMs while infusing them with critical insights aimed at robotic tasks, enhancing tool detection, affordance recognition, and physical concepts understanding. To achieve this, we collect a diverse set of images featuring interactive objects, thus encompassing a broad range of challenges related to object detection, affordance, and physical concept prediction.
@@
Empirical assessments performed in robotic simulators and across various vision task benchmarks substantiate the robust performance of ManipVQA. Our research makes several significant contributions to the fields of robotics and machine learning. Firstly, we propose a novel approach to robotic manipulation and affordance tasks, which addresses the shortcomings of existing methodologies. Secondly, we are dedicated to fostering the research community and have made our datasets, codes, and models publicly available. In addition, this study not only propels the field of robotics forward but also serves as a valuable resource for future investigations in this domain.
Figure 1: The ManipVQA model predictions are in a unified VQA format, which incorporates affordance detection, affordance grounding, and physical concept understanding. If coupled with a heuristic planner, the model can complete manipulation tasks. The masks displayed in the figure are obtained using SAM-HQ.
II Related Works
@@ II-A Large-Object Datasets
The field of computer vision has been significantly advanced by datasets such as ImageNet [8] and PACO [9], which have been crucial for progress in image classification, object detection, and segmentation. While these datasets enhance object instance detection and segmentation and provide a broad semantic understanding of objects, they often overlook the granularity of object parts and attributes, leading to a gap in robotic applications. Addressing this, datasets such as HANDAL [10] and PhysObjects [11] have been introduced, focusing on robotic needs by providing annotations for part masks, object attributes, and affordances, with HANDAL emphasizing manipulable object pose estimation with 308k annotated frames, and PhysObjects offering detailed annotations of household object properties. To unify these advancements, our ManipVQA merges these datasets into a cohesive VQA dataset format, granting models the ability to perform visual reasoning on general images as well as affordance reasoning for manipulation tasks, thereby enhancing the model’s visual understanding for robotic applications.
@@ II-B Multi-Modal Large Language Model
Building on the foundation of extensive large language models like LLaMa [12], which exhibit remarkable language reasoning across various tasks, MLLMs have significantly broadened the scope of language processing by incorporating the ability to understand visual stimuli [13, 2]. A notable innovation in this field is SPHINX [13], which achieves superior performance in multimodal tasks by aligning ensembled visual features with language embeddings through projection layers. Despite these strides, applying MLLMs to areas like robotic manipulation is still in its infancy. Some recent initiatives such as [4, 14, 6] have attempted to integrate robotic-domain knowledge into MLLMs. However, these efforts often overlook critical physical information [4] or do not fully embody a robotic-centric approach [6]. In contrast, our proposed ManipVQA integrates essential physical knowledge and affordance reasoning capabilities from a robotic perspective into MLLMs. We present this information in a direct, human-readable, purely linguistic format, enhancing the utility of MLLMs for practical robotic applications.
@@ II-C Language-Driven Robotics Manipulation
Instruction-based policies in robotics serve as a bridge between high-level human communication and low-level robotic control, offering intuitive interaction and the potential for skill transfer and complex task planning. One approach [15, 16] involves using a pre-trained text encoder to extract text embeddings that encapsulate linguistic knowledge, which are then used to inform robotic policies. For example, CLIPort has obtained the ability to understand semantics and manipulate objects by encoding text input through CLIP[17]. Another research direction leverages the reasoning capabilities of large language models (LLMs) for planning [18, 19] or even directly generating the policy codes [3, 20, 21] by prompt engineering. For example, CaP[22] generates policy codes with detailed comments and context-specific examples to guide LLM output. Furthermore, some studies have trained large-scale policy models [23, 24, 25] using extensive robotic datasets, aiming to learn robotic control policies from diverse data directly. Despite the progress in these areas, a gap remains in the explicit consideration of affordance grounding and physical reasoning, which are essential for proficient manipulation in robotics. Aiming to bridge the gap, our ManipVQA injects robotic affordance and physically grounded information into MLLM while preserving the original reasoning ability.
@@ III Method
Figure 2: An overview of our ManipVQA. We created an extensive vision-language dataset by combining existing resources and expanding affordance grounding tasks using ChatGPT. To maintain consistency with existing VQA datasets, we structured our dataset in a similar VQA format. Utilizing this curated dataset, we then fine-tuned an MLLM. Once integrated with a heuristic policy, the enhanced MLLM is capable of performing a broad array of tasks, notably including complex manipulation tasks.
@@ III-A Modeling of Affordances and Physical Concepts
In the domain of robotic manipulation, understanding and modeling the affordances of objects is critical for enabling robots to interact with their environment effectively. We extend upon the graspable affordance model $\mathbf{A_{grasp}}$ for common tool objects $\mathbf{O_{tool}}$, forming the tuple $(\mathbf{A_{grasp}},\mathbf{O_{tool}})$, as delineated by HANDAL [10]. In our representation, each (partial) object is paired with a bounding box, described by the coordinates $[x_{min},y_{min},x_{max},y_{max}]$, which specify the top-left and bottom-right corners, respectively.
To ensure our model’s adaptability across different contexts, we normalize these coordinates relative to the overall image dimensions, enhancing the model’s generalizability and robustness. Furthermore, we recognize that affordances may vary depending on the specific task $\mathbf{T}$ at hand. For instance, distinct regions of a tool may be utilized for different functions. To address this, we form the tuple $(\mathbf{A_{T}},\mathbf{O_{tool}})$, which associates task-specific affordances $\mathbf{A_{T}}$ with the tool object. The task $\mathbf{T}$ is succinctly described using a brief natural language sentence to encapsulate its functionality.
In addition to affordances, we incorporate the modeling of physical concepts, denoted as $\mathbf{P}_{i}$. These concepts are quantified using discrete levels or boolean values, drawing inspiration from the methodology presented in PhyObjects [11]. Each physical concept $\mathbf{P}_{i}$ is linked to its corresponding object $\mathbf{O}$, resulting in the tuple $(\mathbf{P}_{i},\mathbf{O})$. Table I enumerates the physical concepts under consideration, alongside their respective descriptions. These concepts include but are not limited to, transparency, liquid storage capacity, and sealability—each of which plays a pivotal role in the robot’s ability to interpret and interact with various objects within its environment.
Phy. Concepts Descriptions Type
Transparency Object’s light transmission ability. Levels
Liquids Storage Object’s liquid-holding capacity. Bool
Sealability Object’s closure state. Bool
TABLE I: Physical concepts, descriptions, and representation formats used in ManipVQA.
@@ III-B Instruction Dataset Construction
In the development of the ManipVQA model, we prioritized the assembly of a comprehensive training dataset, amalgamating a variety of publicly available datasets. The aim was to endow the model with a robust understanding of robotic affordance, advanced visual reasoning, and a physically grounded knowledge base.
Robotic Affordance Datasets. The HANDAL dataset [10] serves as a foundation base for imparting the ManipVQA model with the ability to discern objects and identify graspable components. It comprises over 212 real-world objects, each with annotated handle locations. To further refine the model’s capability in understanding complex affordances, we integrated the RGB-D Part Affordance Dataset [26], which differentiates seven pre-defined affordances such as grasp, cut, scoop, contain, pound, support, and wrap-grasp. While the AGD20K dataset [27] represents another extensive affordance resource, it is predominantly labeled with broad human-centric action terms, such as drink or sit-on. Therefore, we have excluded AGD20K from the training regime and instead selected it for zero-shot evaluation, assessing the model’s generalization capabilities.
Physically Grounded Dataset. The PhysObjects dataset [11] was employed to infuse the model with a nuanced understanding of physical properties. This dataset features eight core physical concepts, but we selectively utilized annotations relevant to our focus—specifically, those related to liquid storage suitability, sealability, and transparency as listed in Table I.
General Visual Reasoning Datasets. To maintain and extend the model’s capabilities in general visual reasoning, we incorporated datasets such as the PACO [9], refCoco [28], and the Visual Genome [29], which provide a rich source of information on parts and attributes of common objects,
Augmented Instructions with GPT-4. Considering the limitations of existing annotations in the robotic affordance datasets—primarily confined to partial masks in [10] and basic action types with associated masks in [26], we employed GPT-4 to generate complex and contextually rich affordance-based tasks. This augmentation ensures that our model is not merely trained on overt commands but can also interpret and execute complex, implicitly defined tasks. Listing 1 showcases examples of the prompts used for generating these advanced affordance grounding tasks.
Overall, our meticulously curated training dataset integrates vital annotations that span robotic affordance, physical concept comprehension, and general visual reasoning. The subsequent section will delve into the formulation of robotic tasks that underpin the training of the ManipVQA model.
@@
You are an analytical assistant specializing in
robotic affordance grounding. Your expertise is in
creating tasks that facilitate the training of
robotic policies, enabling robots to reason about
task execution, such as determining the
appropriate part of an object to grasp.
@@
You will be provided with the name of a tool that
can be attached to a robotic arm. The robot is
expected to use this tool to perform a variety of
everyday tasks. Along with the tool name, you will
receive a list of tasks that have already been
generated for this tool.
@@
Diversity: Aim for a wide range of tasks, ensuring
that there is no overlap with previous ones.
@@
Leakage Avoidance: Ensure that the generated tasks
do not explicitly mention the name of the tool object.
@@
with the diversity guideline. Output should be in
the JSON format with the object name as the key.
Listing 1: An example prompt for guiding ChatGPT to generate the affordance grounding task for a specific tool object.
@@
Phys. Gr. Understanding REG-Phy: Liq./Seal./Transp.
@@
Phy 7K
TABLE II: Overview of the ManipVQA Dataset with task template examples. This table summarizes the tasks, their associated capabilities, example templates, and the number of samples in each dataset. The capabilities are represented by the abbreviations: ”Gen.” for General, ”Aff.” for Affordance, and ”Phys. Gr.” for Physically Grounded. The Source Datasets are identified by the acronyms: H for HANDAL [10], P for PACO [9], C for refCoco [28], R for RGB-D Part Affordance Dataset [26], and Phy for PhysObjects [11]. The BBox format is represented by $[x_{min},y_{min},x_{max},y_{max}]$.

### v2

@@ Abstract
While the integration of Multi-modal Large Language Models (MLLMs) with robotic systems has significantly improved robots’ ability to understand and execute natural language instructions, their performance in manipulation tasks remains limited due to a lack of robotics-specific knowledge. Conventional MLLMs are typically trained on generic image-text pairs, leaving them deficient in understanding affordances and physical concepts crucial for manipulation. To address this gap, we propose ManipVQA, a novel framework that infuses MLLMs with manipulation-centric knowledge through a Visual Question-Answering (VQA) format. This approach encompasses tool detection, affordance recognition, and a broader understanding of physical concepts. We curated a diverse dataset of images depicting interactive objects, to challenge robotic understanding in tool detection, affordance prediction, and physical concept comprehension. To effectively integrate this robotics-specific knowledge with the inherent vision-reasoning capabilities of MLLMs, we leverage a unified VQA format and devise a fine-tuning strategy. This strategy preserves the original vision-reasoning abilities while incorporating the newly acquired robotic insights. Empirical evaluations conducted in robotic simulators and across various vision task benchmarks demonstrate the robust performance of ManipVQA. The code and dataset are publicly available at https://github.com/SiyuanHuang95/ManipVQA.
@@ I Introduction
Figure 1: The ManipVQA model generates predictions in a unified VQA format, combining tool-object and affordance detection, affordance grounding, and an understanding of physical object properties. The displayed masks are created using SAM-HQ, based on the bounding boxes predicted by ManipVQA. The contact point, represented by a red star, indicates the geometric center of the bounding box.
@@
Recently, Multi-modal Large Language Models (MLLMs) such as OpenAI’s GPT-4V [1], Google’s Gemini and SPHINX-X [2], have significantly advanced the capabilities for understanding and interpreting both text and images. These models achieve this improvement by aligning multi-modal encoders with Large Language Models (LLMs) through training on numerous text-image pairs or interleaved text-image pairs to enhance a comprehensive understanding of both modalities. As a result, MLLMs exhibit promising potential in addressing common sense reasoning and demonstrate remarkable generalization in vision tasks. However, the application of these models in manipulation tasks [3, 4], such as robotic affordance understanding and the design of associated robotic systems, continues to present significant challenges.
@@
Robotic manipulation [5, 4] is defined as a robot’s capacity to perceive its environment and discern the potential actions applicable to various objects. The direct application of existing MLLMs, pre-trained on common scene reasoning, to robotic manipulation, does not yield satisfactory performance due to the absence of low-level action samples in their pretraining data. Previous research has addressed robotic affordance by prompting MLLMs to process scene images and then generate a sequence of robotic actions, but the performance has been suboptimal. Affordance grounding [6] aims to localize the regions in objects where actions are possible. This task also faces the challenge of establishing an explicit link with object parts due to the diversity of interactive affordances. Prior research [7] employed LLMs to get the answer on affordance with prompt engineering. Despite the potential utility of MLLMs in robotic applications, their effectiveness is limited by several challenges. Classic MLLMs [2] are generally trained on generic image-text pairs, thereby lacking critical robotic knowledge for understanding object affordances and their physical properties. This deficiency in specialized knowledge impedes their performance in manipulation tasks, consequently restricting the types of tasks that robots can execute and the precision of their execution. The gap between the capabilities of MLLMs and the demands of robotic systems constitutes a significant challenge that needs to be addressed.
@@
To address this gap, we present ManipVQA, an innovative framework devised to equip MLLMs with manipulation-centric knowledge via a Visual Question-Answering paradigm. This integration is achieved through a unified VQA approach and a fine-tuning strategy that preserves the original vision-reasoning capabilities of MLLMs while infusing them with critical insights aimed at robotic tasks, enhancing tool detection, affordance recognition, and physical concepts understanding. To achieve this, we collect a diverse set of images featuring interactive objects, thus encompassing a broad range of challenges related to object detection, affordance, and physical concept prediction.
@@
Empirical assessments performed in robotic simulators and across various vision task benchmarks substantiate the robust performance of ManipVQA. Our research makes several significant contributions to the fields of robotics and machine learning. First, we propose a novel approach to robotic manipulation and affordance understanding tasks, which addresses the shortcomings of existing methodologies. Second, we are dedicated to fostering the research community and have made our datasets, codes, and models publicly available.
II Related Work
Figure 2: Overview of ManipVQA: We created a comprehensive vision-language dataset by merging existing datasets and expanding affordance grounding tasks with ChatGPT. To ensure consistency with existing VQA datasets, we adopted a similar VQA format. Using this dataset, we then fine-tuned an MLLM. When combined with a heuristic policy, the enhanced MLLM can perform a wide range of tasks, including complex manipulation tasks.
@@ II-A Large-Object Datasets
The field of computer vision has been significantly advanced by datasets such as ImageNet [8] and PACO [9], which have been crucial for progress in image classification, object detection, and segmentation. While these datasets enhance object instance detection and segmentation and provide a broad semantic understanding of objects, they often overlook the granularity of object parts and attributes, leading to a gap in robotic applications. To address this gap, datasets like HANDAL [10] and PhysObjects [11] have been introduced, focusing on robotic needs by providing annotations for part masks, object attributes, and affordances. HANDAL emphasizes manipulable object pose estimation with 308k annotated frames, while PhysObjects offers detailed annotations of household object properties. To unify these advancements, our ManipVQA merges these datasets into a cohesive VQA dataset format, granting models the ability to perform visual reasoning on general images as well as affordance reasoning for manipulation tasks, thereby enhancing the model’s visual understanding for robotic applications.
@@ II-B Multi-Modal Large Language Model
Building on the foundation of extensive LLMs like LLaMa [12], which exhibit remarkable language reasoning capabilities across various tasks, MLLMs have significantly expanded the scope of language processing by incorporating the ability to understand visual stimuli [13, 2]. A notable innovation in this field is SPHINX [13], which excels in multi-modal tasks by aligning ensembled visual features with language embeddings through projection layers. Despite these advances, the application of MLLMs to robotic manipulation remains in its early stages. Some recent initiatives such as [4, 14, 6] have attempted to integrate robotic-domain knowledge into MLLMs. However, these efforts often overlook critical physical information [4] or do not fully embody a robotic-centric approach [6]. In contrast, our proposed ManipVQA integrates essential physical knowledge and affordance reasoning capabilities from a robotic perspective into MLLMs. We present this information in a direct, human-readable, purely linguistic format, enhancing the utility of MLLMs for practical robotic applications.
@@ II-C Language-Driven Robotics Manipulation
Instruction-based policies in robotics serve as a bridge between high-level human communication and low-level robotic control, offering intuitive interaction and the potential for skill transfer and complex task planning. One approach [15, 16] involves using a pre-trained text encoder to extract text embeddings that encapsulate linguistic knowledge, which are then used to inform robotic policies. For example, CLIPort has obtained the ability to understand semantics and manipulate objects by encoding text input through CLIP[17]. Another research direction leverages the reasoning capabilities of LLMs for planning [18, 19] or even directly generating the policy codes [3, 20, 21] by prompt engineering. For example, CaP[22] generates policy codes with detailed comments and context-specific examples to guide LLM output. Furthermore, some studies have trained large-scale policy models [23, 24, 25] using extensive robotic datasets, aiming to learn robotic control policies from diverse data directly. Despite the progress in these areas, a gap remains in the explicit consideration of affordance grounding and physical reasoning, which are essential for proficient manipulation in robotics. Aiming to bridge this gap, our ManipVQA injects robotic affordance and physically grounded information into MLLM while preserving its original reasoning ability.
@@ III-A Modeling of Affordances and Physical Concepts
In robotic manipulation, understanding and modeling object affordances is essential for enabling robots to interact effectively with their environment. We extend upon the graspable affordance model $\mathbf{A_{grasp}}$ for common tool objects $\mathbf{O_{tool}}$, forming the tuple $(\mathbf{A_{grasp}},\mathbf{O_{tool}})$, as delineated by HANDAL [10]. In our representation, each (partial) object is paired with a bounding box, described by the coordinates $[x_{min},y_{min},x_{max},y_{max}]$, which specify the top-left and bottom-right corners, respectively.
To ensure our model’s adaptability across different contexts, we normalize these coordinates relative to the image dimensions, enhancing the model’s generalizability and robustness. Furthermore, we recognize that affordances may vary depending on the specific task $\mathbf{T}$ at hand. For instance, distinct regions of a tool may be utilized for different functions. To address this, we form a tuple $(\mathbf{A_{T}},\mathbf{O_{tool}})$, which associates task-specific affordances $\mathbf{A_{T}}$ with the tool object. The task $\mathbf{T}$ is succinctly described using a brief natural language sentence to encapsulate its functionality.
In addition to affordances, we incorporate the modeling of physical concepts, denoted as $\mathbf{P}_{i}$. These concepts are quantified using discrete levels or boolean values, drawing inspiration from the methodology presented in PhysObjects [11]. Each physical concept $\mathbf{P}_{i}$ is linked to its corresponding object $\mathbf{O}$, resulting in the tuple $(\mathbf{P}_{i},\mathbf{O})$. Table I enumerates the physical concepts under consideration, alongside their respective descriptions. These concepts include but are not limited to, transparency, liquid storage capacity, and seal-ability. Each concept is crucial for the robot’s ability to perceive and interact with various objects within its environment.
Phys. Concepts Descriptions Type
Transparency Object’s light transmission ability Levels
Liquids Storage Object’s liquid-holding capacity Bool
Seal-ability Object’s closure state Bool
TABLE I: Physical concepts used in ManipVQA.
@@ III-B Instruction Dataset Construction
In the development of the ManipVQA model, we focused on creating a comprehensive training dataset by combining various publicly available datasets. The goal was to equip the model with a strong understanding of robotic affordances, advanced visual reasoning, and a solid foundation in physically grounded knowledge.
III-B1 Robotic Affordance Datasets
The HANDAL Dataset [10] serves as a foundation base for imparting the ManipVQA model with the ability to discern objects and identify graspable components. It comprises over 212 real-world objects, each with annotated handle locations. To further refine the model’s capability in understanding complex affordances, we integrated the RGB-D Part Affordance Dataset [26], which differentiates seven pre-defined affordances such as grasp, cut, scoop, contain, pound, support, and wrap-grasp. While the AGD20K Dataset [27] represents another extensive affordance resource, it is predominantly labeled with broad human-centric action terms, such as drink or sit-on. Therefore, we have excluded AGD20K from the training regime and instead selected it for zero-shot evaluation, assessing the model’s generalization capabilities.
III-B2 Physically Grounded Dataset
The PhysObjects Dataset [11] was employed to infuse the model with a nuanced understanding of physical properties. This dataset features eight core physical concepts, but we selectively utilized annotations relevant to our focus—specifically, those related to liquid storage suitability, seal-ability, and transparency as listed in Table I.
III-B3 General Visual Reasoning Datasets
To maintain and extend the model’s capabilities in general visual reasoning, we incorporated datasets such as the PACO [9], RefCOCO [28], and the Visual Genome [29], which provide a rich source of information on parts and attributes of common objects.
III-B4 Augmented Instructions with GPT-4
Considering the limitations of existing annotations in the robotic affordance datasets—primarily confined to partial masks in [10] and basic action types with associated masks in [26], we employed GPT-4 to generate complex and contextually rich affordance-based tasks. This augmentation enables our model to not only learn explicit commands but also to interpret and execute complex, implicitly defined tasks. Listing 1 showcases examples of the prompts used for generating these advanced affordance grounding tasks.
@@ III-B4 Augmented Instructions with GPT-4
You are an analytical assistant specializing in robotic
affordance grounding. Your expertise is in creating
tasks that facilitate the training of robotic policies,
enabling robots to reason about task execution, such as
determining the appropriate part of an object to grasp.
@@
You will be provided with the name of a tool that can
be attached to a robotic arm. The robot is expected to
use this tool to perform a variety of everyday tasks.
Along with the tool name, you will receive a list of
tasks that have already been generated for this tool.
@@
Diversity: Aim for a wide range of tasks, ensuring that
there is no overlap with previous ones.
@@
Leakage Avoidance: Ensure that the generated tasks do
not explicitly mention the name of the tool object.
@@
with the diversity guideline. Output should be in the
JSON format with the object name as the key.
Listing 1: Example prompt for generating an affordance grounding tasks for a specific tool object using ChatGPT.
@@
Phys. Gr. Understanding REG-Phys: Liq./Seal./Transp.
@@
Phys 7K
TABLE II: Overview of the ManipVQA Dataset: This table summarizes the tasks, their associated capabilities, example templates, and the number of samples in each dataset. The capabilities are represented by the abbreviations: “Gen.” for General, “Aff.” for Affordance, and “Phys. Gr.” for Physically Grounded. The source datasets are identified by the acronyms: H for HANDAL [10], P for PACO [9], C for RefCOCO [28], R for RGB-D Part Affordance Dataset [26], and Phys for PhysObjects [11]. The bounding box (BBox) format is represented by $[x_{min},y_{min},x_{max},y_{max}]$.